Plot the following function on MATLAB

h(t) = cos(pi*t/2) -2 <= t < -1

1 + t3    -1 <= t < 0

1 0 <= t < 1

1 + sin(2*pi*t) 1 <= t < 2

0 elsewhere
Answer #1

If you have any doubt ask in the comment section below.

MATLAB code:

clc; clear all; t1=-2:0.001:-1; h1=cos(pi*t1/2); t2=-1:0.0001:0; h2=1+3*t2; t3=0.001:0.001:1; h3=heaviside(t3); t4=1:0.001:2;

Output:

*2 -1.5 4 -0.5 0 0.5 1 1.5 2